# Adversarial Robustness in Parameter-Space Classifiers

### v1

@@ 4.1 Robustness For 2D Data
Results indicate 3 major trends - First, both MNIST and Fashion-MNIST are adversarially robust. For MNIST, while the TMO and BOTTOM attacks are the only two that managed to incur any above-marginal decrease in classification accuracy, it is still far below the accuracy decrease of around $60\%$ that is traditionally observed in similarly-bounded PGD attacks over signal-domain classifiers without any robust training ((Wong et al., 2020)). While higher-bounded attacks over the Fashion-MNIST dataset have been more successful, we claim their success is substantially lower than the effect of a similarly constrained attack over a signal-domain classifier. To show this, we train four baseline signal-domain classifiers and attack with PGD using similar $\mathcal{L}_{\infty}$ norm constraints and PGD iteration number. We use Vision-Transformers (ViTs) (Dosovitskiy, 2020) due to their outstanding record in natural image classification (Han et al., 2022), and two convolutional models for their ubiquity and foundational role in computer-vision. To experiment with models of different sizes we train a ViT-Large model and a Resnet18, as well as a ViT-Base model and a compact,2-layer CNN. Full architecture details are included in Section A.3.2.
We compare the robust accuracy of these four Signal-space classifiers under standard 100-iteration PGD attack, with that of our parameter-space classifier subjected to all attacks from Sections 3.3.1, 3.3.2, 3.3.3, 3.3.4 and 3.3.5. Results, shown in figure 3, show parameter-space classifiers show significantly higher robustness, especially for lower attack bounds.
The second indicated trend is that we observe that the full PGD attack demonstrates lower adversarial prowess compared to BOTTOM and TMO attacks in all experiments. We attribute this to the fact that in full-PGD the numerous steps of second-order optimization required for standard PGD in our case encourage vanishing gradients phenomena, namely attenuating gradient information backpropagated through the modulation optimization process $R$. These difficulties are added to the computational encumbrance of running full PGD (see section 4.4), and further stress the need in initial methods developed here (sections 3.3.2&3.3.3), as well as further related research. While computationally lightest, implicit differentiation proves less potent than most explicit methods. We attribute this to the strong dependence of this method in fulfillment of specified optimality conditions.
Lastly, we conclude that the INR-domain attack (ICOP) failed to impose any actual effect on classification accuracy. We attribute this result to the heuristic nature of the approaches used to constrain the INR-domain perturbation’s response in the signal domain (section 3.3.4). We still include it within our suite of proposed attacks, as it proved more efficient when initiated over 3D data (section 4.2).
Figure 3: Robust accuracy under $\mathcal{L}_{\infty}$ norm constraints - for parameter-space (solid curves) and signal-space (dotted curves) classifier attacks.
@@ 4.2 Robustness For 3D Data
In section 4.1 we survey adversarial robustness of classifiers INR for 2D data. The importance of this baseline is comparability to the widely studied adversarial robustness of 2D classifiers. Nonetheless, practical advantages of using INRs for 2D data are usually limited, since common grid representations suffice for most common tasks. Data modalities of larger dimensions, such as 3D data, however, more commonly gain from using INRs for achieving downstream tasks, including classification ((Jiang et al., 2020; Molaei et al., 2023; Niemeyer et al., 2020)). We use the ModelNet 10 dataset ((Wu et al., 2015)) consisting of 4900 3D CAD models given in Mesh-grid representation.
@@ 4.2.1 Signal Space Attack
To demonstrate inherent adversarial robustness over 3D data under the BVA attack (section 3.4), we must first demonstrate its prowess in attenuating signal-domain classification accuracy. To do so we train a voxel-grid classification model based on Point-Voxel CNN layers proposed by (Liu et al., 2019b). We train the classifier over ModelNet10, converting every mesh-grid sample to a voxel-grid of dimension $15\times 15\times 15$ using Trimesh ((Dawson-Haggerty et al., )). Then, we attack this trained model with BVA for a wide range of $\mathcal{L}_{0}$ bounds and assess attacked classification accuracy. Results are displayed in figure 4, showing BVA is efficient over the tested signal-space classifier, inducing exponential decay of classification accuracy with growth in number of flipped voxel bits.
Figure 4: BVA robust accuracy for signal-space classifier and weight-space classifier - under $\mathcal{L}_{0}$ norm constraints.
@@ 4.2.2 Parameter Space Attack
Figure 4 compares the robust accuracy of the signal domain (black) and parameter-space classifiers, where the parameter-space classifier is attacked with TMO, BOTTOM, ICOP, and implicit attacks. We note that we do not experiment with the full PGD attack in parameter-space for the 3D INR classifier, since to optimize INR for the ModelNet10 dataset with sufficient input signal fidelity, we had to perform 500 modulation optimization iterations per sample. While computation times for doing so are not very high (several seconds), backpropagation through these 500 optimization steps calls for infeasible memory and compute time requirements, and also lead to severe vanishing/exploding gradient issues. We further elaborate on these considerations in Section 4.4.
While the parameter-space classifier exhibits lower clean accuracy (73% in parameter-space against 82% in signal space), results demonstrate conclusive increased adversarial robustness for the signal-space classifier. We can specify two major trends - the first is that the absolute decrease in classification accuracy is much lower for the parameter-space classifier. The second is that for the attacked weight-space classifier, the decrease in accuracy attenuation across bound increase saturates significantly faster. The data from Figure 4 is presented numerically in Table 2.
@@ 4.3 Qualitative Results
In this section we provide qualitative evidence for the inherent robustness of parameter-space classifiers. Adversarial attacks are known to effectively exploit the amplifications of small input-space perturbations (Goodfellow, 2014; Haleta et al., 2021). As previously discussed, we argue that the enhanced robustness of parameter-space classifiers arises from the attacker’s inability to directly perturb the classifier’s input. Instead, the attacker must optimize a constrained signal-domain adversarial perturbation for which the response received as input to the classifier (after modulation optimization Equation 1) is adversarially potent. This is deemed to be substantially more challenging for the attacker, as modulation optimization tends to attenuate adversarial prowess fed as input to the classifier.
To illustrate this phenomenon, we report the mean amplification, measured as the difference in activation values between perturbed and clean input data, for each layer in our pipeline (Figure 2). These measurements, shown in Figure 5, were calculated using the MNIST dataset, with every sample propagated through the pipeline after being perturbed by each attack in our suite with a constraint of $\varepsilon=\frac{32}{256}$. To further validate the relationship between attack success and amplification, we differentiate between robust and non-robust samples. Successful attacks demonstrate an average amplification approximately three times greater than that of failed attacks. In both cases, amplification sharply plummets upon entering the first classifier layer (layer 10) and remains consistently low in subsequent layers.
These findings indicate that adversarial amplification within the INR layers has a negligible impact on the downstream classification model. This supports our claim that modulation optimization effectively obfuscates the adversarial potency of perturbations applied in the signal domain. Importantly, this effect is not observed for signal-domain classifiers, as we show in Figure 6. Results for other attack constraints and datasets are reported in Appendix B.
@@
In Figure 1 we demonstrate the affect of adversarial perturbations applied in signal-domain over the latent representations attained from perturbed data, and compare between parameter-space classifiers and traditional signal-space classifiers. To this end, as a signal-domain baseline, we first train a convolutional autoencoder over the MNIST dataset. For a fair comparison, this autoencoder is trained disjointly from the downstream classification task. We then train a classifier over the learned latent space, and finally perform 100 iterations of signal-domain adversarial attack over this classifier. The bottom half of Figure 1 illustrates t-SNE projection of the trained autoencoder’s latent representations before (left) and after (right) the adversarial attack. The top half of the plot corresponds to parameter-space classifiers, showing t-SNE projections of INR modulations optimized for clean (right) data and for data attacked with our full-PGD attack (Section 3.3.1. Prior to the attacks, the latent space of both signal and parameter-space domains sis well-structured, with clear separation between the different classes. However, after the attack, the model operating strictly in signal-domain (i.e. the autoencoder baseline) experiences severe disruption of this structure – classes of adversarial samples within the classifier’s latent input space are no longer distinct and are instead intermixed. This indicates a significant deterioration in the class separability, highlighting the impact of the adversarial perturbations on the latent space. Adversarial attacks optimized for the parameter-space classifier, on the other hand, seem inconsequential to the input-space structure of the parameter-space classifier, including class separability. These conclusions are well-correlated with those from Figure 5.
4.4 Computational Resource Analysis
In sections 4.1,4.2 we provide evidence for the adversarial robustness of parameter-space classifiers based on classification accuracies under the presence of adversarial perturbations. In this section, we further support these claims from a different, practical perspective. We show that orthogonally to the efficacy of adversarial attacks over parameter-space classifiers, attacking these classifiers also poses significant added computational difficulties for the potential attacker. These difficulties stem from the optimization process inherent within the classification pipeline’s inference process. To illustrate this, we measure the computation times required for a single modulation optimization step during a single PGD iteration on an NVIDIA RTX-2080 GPU. Our results show that for the MNIST dataset with 512-dimensional modulation vectors, the compute time per step is approximately 0.015 seconds, while for the ModelNet10 dataset with 2048-dimensional modulation vectors, it is 0.02 seconds. Even for a small number of PGD steps, increasing the number of modulation optimization steps – an aspect controlled by the model, not the attacker – increases attack optimization times within several orders of magnitude. Importantly, this increase does not significantly impact the computation time required for ”clean” (unperturbed) inference. For instance, with 100 modulation optimization steps set by the model to solve Equation 1, ”clean” inference would require 1.5 seconds per sample, while perturbation optimization of 100 PGD iterations would require the attacker to invest 150 seconds per-sample.
5 Discussion
5.1 Conclusion
In this work, we’ve explored the adversarial robustness of parameter-space classifiers. We’ve presented the challenges involved in translating common PGD-based attacks onto models performing classification directly in deep parameter-space, and addressed these challenges by developing a novel set of five different adversarial attacks designed for parameter-space classification, applying perturbations both in the signal domain and in parameter space. We then tested these attack over three different datasets from two popular modalities, and shown that parameter-space classifiers exhibit an increased inherent adversarial robustness, compared to their counterparts classifying in signal-domain. We’ve further presented the notion of ”robustness” stemming from the potentially large added computational resources required for certain cases of parameter-space classifier attacks.
5.2 Limitations and Future Work
While we consider our work an important initial study of the currently under-explored topic of adversarial robustness in parameter-space classification, several notable limitations remain. First, in this work we consider a relatively limited set of simple datasets from both 2D and 3D modalities. This choice mainly stems from the fact that current state-of-the-art parameter space classifiers have yet to achieve adequate ”clean” (non-attacked) classification results over more challenging datasets ((Navon et al., 2023a, b)). Since the focus of our work is adversarial robustness of parameter-space classifiers, experimenting with our proposed attacks over classifiers with low ”clean” accuracy would make assessment of our work less reliable. We hope future exploration of methods for better non-adversarial parameter-space classification would expand our set of potential datasets for robustness exploration. Second, we note that the adversarial robustness potentially attained when using parameter-space classifiers often comes with the cost of reduced classification accuracy, as shown in section 4. Finally, our work focuses on inherent adversarial robustness, and does not discuss active defenses. We deem our findings about inherent robustness an encouraging starting point for future research in development of adversarial defenses and robust modeling.
@@ A.1.1 MNIST & Fashion-MNIST
All experiments described in this section have been conducted on a single NVIDIA RTX-2080 GPU. Following (Dupont et al., 2022), for each of the three datasets considered in this paper, we first train a SIREN (Sitzmann et al., 2020) jointly over the entire training set. For MNIST and Fashion-MNIST we optimize modulation vectors of dimension $512$ for 6 epochs of 3 modulation optimization steps per-sample, with a learning-rate of $0.000005$ and a batch size of $128$. We use 10 SIREN layers with hidden dimension of $256$. We then proceed to creating the Functaset , which is the dataset of modulation vectors conditioning the SIREN for per-sample representation. For each sample we optimize for 10 modulation optimization steps, with a learning-rate of 0.01.
@@ A.1.2 ModelNet10
For ModelNet10 we similarly follow (Dupont et al., 2022), however we first convert the Mesh-grid data onto a voxelgrid, as previously mentioned (Section 3.4). Each voxelgrid is resampled to dimension $15\times 15\times 15$. Due to the increased dimensionality, we train $2048$-dimensional Functa modulations with a SIREN similar to the one used for the MNIST dataset. After training the shared SIREN weights, each modulation vector is further optimized for $500$ modulation optimization steps.
@@ A.2 Adversarial Attacks
A.3 Baseline Models
A.3.1 Classifier Baselines
In this section we specify implementation and training details related to the signal-space classifier baseline CNN and ViT architectures used in Section 4. In section 4.1 we train 4 signal-domain classifiers – a small CNN, a ResNet18, ViT-Large/16 and Vit-Base/16. For the ViTs we leverage pre-trained models from (Wightman, 2019), and finetune for our classification task using a a linear probe prior to a applying Softmax. For the Resnet we train from scratch, and project encoded data with a linear probe followed by a Softmax. For the small-CNN we use two convolutional layers with ReLU activations, and a single MaxPooling operation between them. For all baselines we train over the Fashion-MNIST training set for 5 epochs with learning-rate of 0.0001.
A.3.2 Autoencoder Baselines
In Section 4.3, we present a qualitative analysis of adversarial perturbation amplification across pipeline layers, along with t-SNE projections of modulation spaces before and after adversarial perturbation. To connect these observations to the inherent adversarial robustness of parameter-space classifiers, we include a comparison with a signal-space classifier. In order to propose a signal-space classifier with an appropriate counterpart to the explored parameter-space model modulation space, we choose to train classifiers operating over latent spaces trained by autoencoders. We model each autoencoder as a 3-layer CNN with ReLU activations and train with a self-supervised reconstruction objective for 10 epochs. After the autoencoder is trained, we freeze it and train a linear classifier with 2 hidden layers and a single ReLU activation on top of it. The frozen encoder layers acts as the INR counterpart in Figure 5, and the MLP head is parallelized to the parameter-space classifier. We use a latent dimension of $128$ for all encoders.

### v2

@@ 4.1 Robustness For 2D Data
Results indicate 3 major trends - First, both MNIST and Fashion-MNIST are adversarially robust. For MNIST, while the TMO and BOTTOM attacks are the only two that managed to incur any above-marginal decrease in classification accuracy, it is still far below the accuracy decrease of around $60\%$ that is traditionally observed in similarly-bounded PGD attacks over signal-domain classifiers without any robust training (Wong et al. (2020)). While higher-bounded attacks over the Fashion-MNIST dataset have been more successful, their success is substantially lower than the effect of a similarly constrained attack over a signal-domain classifier. To show this, we train four baseline signal-domain classifiers and attack with PGD using similar $\mathcal{L}_{\infty}$ norm constraints and PGD iteration number. For a representative pool of common signal-domain classifiers we use ViT-Base/16, ViT-Large/16 (Dosovitskiy (2020)), a Resnet18 and a smaller 2-layer CNN. Full architecture details are in section A.4.2. We compare the robust accuracy of these four Signal-space classifiers under standard 100-iteration PGD attack, with that of our parameter-space classifier subjected to all attacks from sections 3.3.1, 3.3.2, 3.3.3, 3.3.4 and 3.3.5. Results, shown in figure 2(a), indicate parameter-space classifiers show significantly higher robustness, especially for lower attack bounds.
Furthermore, we observe that the full PGD attack demonstrates lower adversarial prowess compared to BOTTOM and TMO attacks in all experiments. We attribute this to the fact that in full-PGD the numerous steps of second-order optimization required for standard PGD in our case encourage vanishing gradients phenomena, namely attenuating gradient information backpropagated through the modulation optimization process $R$. While computationally lightest, implicit differentiation proves less potent than most explicit methods. We attribute this to the strong dependence of this method in fulfillment of specified optimality conditions.
(a) Fashion-MNIST robust accuracy in $\mathcal{L}_{\infty}$.
(b) ModelNet10 robust accuracy in $\mathcal{L}_{0}$.
Figure 2: Robust accuracy for parameter-space (solid curves) and signal-space (dotted curves) classifiers.
@@ 4.2 Robustness For 3D Data
In section 4.1 we survey adversarial robustness of classifiers INR for 2D data. The importance of this baseline is comparability to the widely studied adversarial robustness of 2D classifiers. Nonetheless, practical advantages of using INRs for 2D data are usually limited, since common grid representations suffice for most common tasks. Data modalities of larger dimensions, such as 3D data, however, more commonly gain from using INRs for achieving downstream tasks, including classification (Jiang et al. (2020); Molaei et al. (2023); Niemeyer et al. (2020)). In our experiments we use the ModelNet 10 dataset (Wu et al. (2015)) consisting of 4900 3D CAD models given in Mesh-grid representation.
@@ 4.2.1 Signal Space Attack
To demonstrate inherent adversarial robustness over 3D data under the BVA attack (section 3.4), we must first demonstrate its prowess in attenuating signal-domain classification accuracy. To do so we train a voxel-grid classification model based on Point-Voxel CNN layers proposed by Liu et al. (2019b). We train the classifier over ModelNet10 (full optimization details brought in section A.1.2) and then attack this trained model with BVA for a wide range of $\mathcal{L}_{0}$ bounds and assess attacked classification accuracy. Results are displayed in figure 2(b), showing BVA is efficient over the tested signal-space classifier, inducing exponential decay of classification accuracy with growth in number of flipped voxel bits.
@@ 4.2.2 Parameter Space Attack
Figure 2(b) compares the robust accuracy of the signal domain (black) and parameter-space classifiers, where the parameter-space classifier is attacked with TMO, BOTTOM, ICOP, and implicit attacks. We note that we do not experiment with the full PGD attack in parameter-space for the 3D INR classifier, since to optimize INR for the ModelNet10 dataset with sufficient input signal fidelity, we found it necessary to perform 500 modulation optimization iterations per sample. While computation times for doing so are not very high (several seconds), backpropagation through these 500 optimization steps calls for infeasible memory and compute time requirements, and also lead to severe vanishing/exploding gradient issues. We further elaborate on these considerations in appendix C.
While the parameter-space classifier exhibits lower clean accuracy (73% in parameter-space against 82% in signal space), results demonstrate conclusive increased adversarial robustness for the signal-space classifier. We can specify two major trends - the first is that the absolute decrease in classification accuracy is much lower for the parameter-space classifier. The second is that for the attacked weight-space classifier, the decrease in accuracy attenuation across bound increase saturates significantly faster. The data from fig. 2(b) is presented numerically in table 2.
@@ 4.3 Qualitative Results
In this section we provide qualitative evidence for the inherent robustness of parameter-space classifiers. As previously discussed, we argue that the enhanced robustness of parameter-space classifiers arises from the attacker’s inability to directly perturb the classifier’s input. Instead, the attacker must optimize a constrained signal-domain adversarial perturbation for which the response received as input to the classifier (after modulation optimization eq. 1) is adversarially potent. This is deemed to be substantially more challenging for the attacker, as modulation optimization tends to attenuate adversarial prowess fed as input to the classifier. To illustrate this phenomenon, we report the mean amplification, measured as the difference in activation values between perturbed and clean input data, for each layer in our pipeline (fig. 1). These measurements, shown in fig. 3, were calculated for classification over the MNIST dataset, with every sample propagated through the pipeline after being perturbed by each attack in our suite with a constraint of $\varepsilon=\frac{32}{256}$. To further validate the relationship between attack success and amplification, we differentiate between robust and non-robust samples. Successful attacks demonstrate an average amplification approximately three times greater than that of failed attacks. In both cases, amplification sharply plummets upon entering the first classifier layer (layer 10) and remains consistently low in subsequent layers. These findings indicate that adversarial amplification within the INR layers has a negligible impact on the downstream classification model. This supports our claim that modulation optimization effectively obfuscates the adversarial potency of perturbations applied in the signal domain. Importantly, this effect is not observed for signal-domain classifiers, as we show in fig. 5. Results for other attack constraints and datasets are reported in appendix B.
@@
In fig. 4 we demonstrate the affect of adversarial perturbations applied in signal-domain over the latent representations attained from perturbed data, and compare between parameter-space classifiers and traditional signal-space classifiers. To this end, as a signal-domain baseline, we first train a convolutional autoencoder over the MNIST dataset (section A.4.2), disjointly from the downstream classification task. We then train a classifier over the learned latent space, and finally perform 100 iterations of signal-domain adversarial attack over this classifier. The bottom half of fig. 4 illustrates t-SNE projection of the trained autoencoder’s latent representations before (left) and after (right) the adversarial attack. The top half of the plot corresponds to parameter-space classifiers, showing t-SNE projections of INR modulations optimized for clean (right) data and for data attacked with our full-PGD attack (section 3.3.1. Prior to the attacks, the latent space of both signal and parameter-space domains is well-structured, with clear separation between the different classes. However, after the attack, the model operating strictly in signal-domain (i.e. the autoencoder baseline) experiences severe disruption of this structure – classes of adversarial samples within the classifier’s latent input space are no longer distinct and are instead intermixed. This indicates a significant deterioration in the class separability, highlighting the impact of the adversarial perturbations on the latent space. Adversarial attacks optimized for the parameter-space classifier, on the other hand, seem inconsequential to the input-space structure of the parameter-space classifier, including class separability. These conclusions are well-correlated with those from fig. 3.
Figure 4: t-SNE projection of modulation vectors fitted for clean and adversarially-perturbed data - for parameter-space and signal-space classifiers.
5 Conclusion
In this work, we investigate the adversarial robustness of parameter-space classifiers. We highlight the challenges in adapting commonly used PGD-based attacks to models that perform classification directly in the deep parameter space. To address these challenges, we propose a novel set of five distinct adversarial attacks tailored for parameter-space classification. Our results demonstrate that parameter-space classifiers exhibit enhanced inherent adversarial robustness compared to those operating strictly in the signal domain. While this study serves as a significant initial exploration of adversarial robustness in parameter-space classification—a relatively under-explored area—we acknowledge several limitations. First, current state-of-the-art parameter-space classifiers still struggle to achieve satisfactory ”clean” (non-attacked) classification performance on more complex datasets (Navon et al. (2023a; b)). Consequently, we focus on a limited set of relatively simple datasets from both 2D and 3D modalities, where ”clean” accuracy is sufficient for a reliable evaluation of adversarial attack effects. We hope that future advancements in non-adversarial parameter-space classification will broaden the range of datasets available for robustness studies. Second, we note that the adversarial robustness observed in parameter-space classifiers often comes at the expense of reduced classification accuracy, as demonstrated in section 4.
@@ A.1.1 MNIST & Fashion-MNIST
All experiments described in this section have been conducted on a single NVIDIA RTX-2080 GPU. Following Dupont et al. (2022), for each of the three datasets considered in this paper, we first train a SIREN Sitzmann et al. (2020) jointly over the entire training set. For MNIST and Fashion-MNIST we optimize modulation vectors of dimension $512$ for 6 epochs of 3 modulation optimization steps per-sample, with a learning-rate of $0.000005$ and a batch size of $128$. We use 10 SIREN layers with hidden dimension of $256$. We then proceed to creating the Functaset , which is the dataset of modulation vectors conditioning the SIREN for per-sample representation. For each sample we optimize for 10 modulation optimization steps, with a learning-rate of 0.01.
@@ A.1.2 ModelNet10
For ModelNet10 we similarly follow Dupont et al. (2022), however we first convert the Mesh-grid data onto a voxelgrid, as previously mentioned (section 3.4). Each voxelgrid is resampled to dimension $15\times 15\times 15$. Due to the increased dimensionality, we train $2048$-dimensional Functa modulations with a SIREN similar to the one used for the MNIST dataset. After training the shared SIREN weights, each modulation vector is further optimized for $500$ modulation optimization steps.
@@ A.2 Adversarial Attacks
A.3 Binary Optimization in BVA
In BVA eq. 5 we formulate the attacker objective as:
Since voxel-grid data is represented by binary mask, our optimized perturbation $\delta$ in this case is a binary mask governing the bit-flip imposed over the original, ”clean” voxel-grid data. This formulation calls for the incorporation of binary optimization of the mask $\delta$. Trivial binarization of $\delta$ (e.g. by thresholding) is non-differentiable. To optimize $\delta$ as a binary mask we follow the binary optimization technique presented in Iliadis et al. (2020). This method maintains two congruent masks - a binary mask $\Phi$ and a real mask $\Phi_{c}$. $\Phi$ is used for forward and backpropagation, where each gradient step applies the gradients calculated for $\Phi$ over $\Phi_{c}$. The latter is then binarized (without gradient tracking) to produce $\Phi$ for the next iteration.
In our experiments, we optimize under varying $\mathcal{L}_{0}$ constraints (fig. 2(b)), using a learning-rate of $0.01$.
A.4 Baseline Models
A.4.1 Classifier Baselines
In this section we specify implementation and training details related to the signal-space classifier baseline CNN and ViT architectures used in section 4. In section 4.1 we train 4 signal-domain classifiers – a small CNN, a ResNet18, ViT-Large/16 and Vit-Base/16. For the ViTs we leverage pre-trained models from Wightman (2019), and finetune for our classification task using a a linear probe prior to a applying Softmax. For the Resnet we train from scratch, and project encoded data with a linear probe followed by a Softmax. For the small-CNN we use two convolutional layers with ReLU activations, and a single MaxPooling operation between them. For all baselines we train over the Fashion-MNIST training set for 5 epochs with learning-rate of 0.0001.
A.4.2 Autoencoder Baselines
In section 4.3, we present a qualitative analysis of adversarial perturbation amplification across pipeline layers, along with t-SNE projections of modulation spaces before and after adversarial perturbation. To connect these observations to the inherent adversarial robustness of parameter-space classifiers, we include a comparison with a signal-space classifier. In order to propose a signal-space classifier with an appropriate counterpart to the explored parameter-space model modulation space, we choose to train classifiers operating over latent spaces trained by autoencoders. We model each autoencoder as a 3-layer CNN with ReLU activations and train with a self-supervised reconstruction objective for 10 epochs. After the autoencoder is trained, we freeze it and train a linear classifier with 2 hidden layers and a single ReLU activation on top of it. The frozen encoder layers acts as the INR counterpart in fig. 3, and the MLP head is parallelized to the parameter-space classifier. We use a latent dimension of $128$ for all encoders.